40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 11 00:37:21 2013, Masayuki, Update, Green Locking, Y-arm ALS  OLTF.pdfERR.pdf
    Reply  Tue Oct 15 01:39:07 2013, Masayuki, Update, Green Locking, Y-arm ALS  PDHinloop.pdfALSinloop.pdf
Message ID: 9232     Entry time: Fri Oct 11 00:37:21 2013     Reply to this: 9240
Author: Masayuki 
Type: Update 
Category: Green Locking 
Subject: Y-arm ALS  

[Manasa, Masayuki]

- Motivation

Our goal is to realize PRMI+one arm again. However we found that the noise level of the Y-arm is worse than before (entry).
  Today we went through into the servo gains of the ALS related loops. 

- What we did

Step 1 to 6 is for Yarm
Alignment of the cavity and the green:
1. Locked arms using IR PDH, aligned the green beam to increase the transmission. Now the value of ALS-TRY_OUTPUT is more than 0.8.

Checking and adjustment of the end green PDH gain:
2. Measured the OLTF of green PDH loop.
3. The gain of the PDH box was 8.2. We found that the UGF was too high and the phase mergin was too low (20deg)
    Therefore, the gain was reduced to the gain to 6.8. Now, the UGF and phase margin are 17.7 kHz, 41.96 degree, respectively.

Phase tracker loop:
4. Measured the OLTF of the phase tracker loop. The UGF was 2 kHz, and phase margin was 45 degree.
    We found that these were already the nominal and optimized numbers.
    For a reference: the filter bank C1:ALS-BEATX_FINE_PHASE has the gain of 110.

ALS loop:
5. Disable the IR PDH lock, and stabilized Yarm by ALS. We measured the OLTF of the ALS loop (attachment 1).
    The UGF and phase margin were turned out to be 125 Hz and 41 degree. respectively. This looks pretty optimal.
    The ALS servo gain (the gain of the C1:ALS-YARM module) was 15.0.
6. We measured the in-loop noise of the ALS loop (C1:BEATY_FINE_PHASE_OUT_HZ) (attachment 2).
    The comparison of the in-loop performance is discussed below.

- Discussion

After these adjustment, we found that the ALS in-loop noise of Yarm decreased in high frequency band.
(see this entry for the comparison. Sorry for my laziness! I don't have the overlaid plot)

If we believe this is true, lowering the end PDH gain improved the noise level between 100Hz to 1kHz.
This sounds weird as we decreased the PDH gain, rather than increased. We should confirm this effect by increasing the gain.

Now the in-loop RMS is started to be dominated by the peaks at 3, 16, and 24 Hz.
We should compare the current in-loop spectrum with the previous spectrum when the ALS was working fine.

By the way:
We suffered from frequent disruptions of the ALS servo during our investigation.
As we speculated that this was caused by the malfunction of the green PDH loop, we left the arm still and observed
how the green PDH lock is robust. Our discovery was that the green PDH loop had frequent interruptions (every 5~10min).

From this observation, we strongly feel that we need to look into the entire end PDH loop.

Attachment 1: OLTF.pdf  8 kB  | Hide | Hide all | Show all
OLTF.pdf
Attachment 2: ERR.pdf  27 kB  | Show | Hide all | Show all
ELOG V3.1.3-